Can you please help me with this equation? 2x-1= x^2-4x+4
Answer by ankor@dixie-net.com(22740)

2x-1= x^2-4x+4
:
Write it:
x^2 - 4x + 4 = 2x - 1
:
Subtract 2x and add 1 to both sides:
x^2 - 4x - 2x + 4 + 1 = 0
:
x^2 - 6x + 5 = 0; a quadratic equation
:
Factor this to:
(x - 5)(x - 1) = 0
:
x = +5
and
x = +1
:
Confirm this by substituting 5 for x in the original equation
2x - 1 = x^2 - 4x + 4
2(5) - 1 = 5^2 - 4(5) + 4
10 - 1 = 25 - 20 + 4
9 = 5 + 4;
:
You can do the same with x = 1